# LLM as OS (llmao), Agents as Apps: Envisioning AIOS, Agents and the AIOS-Agent Ecosystem

### 06 Dec 2023 v1

@@ Abstract
In this paper, we begin by introducing the architecture and historical evolution of traditional Operating Systems (OS). Then we formalize a conceptual framework for AIOS through “LLM as OS (LLMAO),” drawing analogies between AIOS components and traditional OS elements: LLM is likened to OS kernel, context window to memory, external storage to file system, hardware tools to peripheral devices, software tools to programming libraries, and user prompts to user commands. Subsequently, we introduce the new AIOS-Agent Ecosystem, where users and developers can easily program Agent Applications (AAPs) using natural language, democratizing the development of and the access to computer software, which is different from the traditional OS-APP ecosystem, where desktop or mobile applications (APPs) have to be programmed by well-trained software developers using professional programming languages. Following this, we explore the diverse scope of Agent Applications. These agents can autonomously perform diverse tasks, showcasing intelligent task-solving ability in various scenarios. We delve into both single agent systems and multi-agent systems, as well as human-agent interaction. Lastly, we posit that the AIOS-Agent ecosystem can gain invaluable insights from the development trajectory of the traditional OS-APP ecosystem. Drawing on these insights, we propose a strategic roadmap for the evolution of the AIOS-Agent ecosystem. This roadmap is designed to guide the future research and development, suggesting systematic progresses of AIOS and its Agent applications.
@@ 1 Introduction
(b) Architecture of Large Language Model as OS (LLMAO) for AIOS.
Figure 2: Illustrations of the architectures of OS and AIOS (LLMAO).
@@
Inspired by the architecture of traditional OS (as shown in Figure. 1(a)), we present a general framework for LLM as OS (LLMAO) with several key components in Section 3 (as shown in Figure. 1(b)): LLM as Kernel, Context Window as Memory, External Storage as File, Tools as Devices/Libraries, User Prompt/Instruction as User Interface (UI), and Agents as Applications. The conceptual framework presented draws an analogy between an LLM as OS (LLMAO) and a traditional Operating System (OS), mapping various components of the LLMAO to elements of an OS. The LLM itself is likened to the kernel, the central part managing the system’s core functions. The LLM’s context window is compared to the memory of an OS, handling immediate context selection and data processing (Shi et al., 2023). External storage for the LLM is analogous to the files of an OS, allowing for long-term data storage. Meanwhile, there are corresponding data retrieval methods to enable retrieval-augmented LLMs (Guu et al., 2020), which serve as the file system of OS to manage and find relevant files. Besides, LLM can make use of various tools for task solving (Ge et al., 2023), including both hardware tools and software tools. Hardware tools in the LLMAO framework are equated to peripheral devices in a traditional OS, each offering specific functionalities to help interact with the physical world, while software tools in the LLMAO framework serve as the programming libraries in traditional OS, enabling Agent applications to interact with the virtual/digital world. User prompts or instructions for the LLM are akin to the user interface (UI) in a traditional OS, facilitating interaction between the user and the system. The user prompts or instructions can be direct natural language instructions provided by the user, and they may also be (sometimes semi-structured) natural language instructions converted from users’ non-natural language instructions such as clicking on icons. This mapping provides a systematic way to understand the operational similarities between LLMAO-based AIOS and traditional OS.
@@
Upon establishing a robust conceptual framework for LLMAO-based AIOS, we introduce the AIOS-Agent Ecosystem, akin to the traditional OS-APP Ecosystem, in Section 4. We begin by introducing the concept of Agent Applications (AAPs) within LLMAO, analogous to traditional applications (APPs) based on an operating system. These AAPs represent a diverse scope of specialized tasks for users to execute based on LLMAO. As illustrated in Figure 4, by integrating the LLMAO layer with the OS layer, Hardware layer, and the Agent Application layer, we can construct an autonomous AI Agent system. Such AI Agent system responds to user prompts or instructions in natural language and is capable of performing a multitude of tasks based on its interaction with the physical or digital environment. Since a diverse scope of Agents can be developed on top of the shared AIOS foundation, this eventually leads to an AIOS-Agent ecosystem. Moreover, in the new AIOS-Agent Ecosystem, the users and developers can easily program Agent Applications (AAPs) using natural language, democratizing the development of and the access to computer software, which is different from the traditional OS-APP ecosystem, where desktop or mobile applications (APPs) have to be programmed by well-trained software developers using professional programming languages.
@@
Following this, we further delve into the practical LLMAO-based Agent Applications in Section 5. This section explores the potential of enhancing LLMAO’s functionality by developing various agents in the real world, and further investigates the dynamic interactions between multiple agents and humans. LLMAO-based agents are characterized by their creative autonomy, enabling them to generate novel ideas, narratives, or solutions not pre-programmed into them (Chase, 2022; Gravitas, 2023; Ge et al., 2023; Li et al., 2023; Yao et al., 2022b, a), which is indicative of an advanced level of creative intelligence (Yuan et al., 2022; Franceschelli and Musolesi, 2023). Specifically, Section 5.1 discusses applications of single agents, Section 5.2 explores multi-agent systems, and Section 5.3 focuses on human-agent interactions.
@@
Finally, in Section 6, we explore several crucial future research directions for LLMAO and AIOS in general, drawing parallels and learning from the evolution of traditional operating systems. These directions span a wide array of areas, aiming to enhance the capabilities and applications of LLMAO: 1) Resource Management. For example, OSes employ virtual and shared memories to address the issue of limited physical memory. LLMAOs can inspire from these ideas to mitigate their own problem of limited context window challenges; 2) Communication. Different OSes and applications communicate using standardized protocols (such as Domain-Specific Languages); LLMAOs and Agents can build and use similar standard protocols for exchanging data and instructions with various systems, ensuring compatibility and smooth interaction across diverse platforms; 3) Security. Security vulnerabilities in OSes are important issues. State-of-the-art approaches aim to detect and capture malware and viruses at various levels. Similarly, LLMAOs can implement detection and intervention mechanisms to regulate and monitor the implementation of third-party tools and Agent Applications.
@@ 2.1.3 Operating System Ecosystem
Similarly, we envision an AIOS-Agent ecosystem, where LLM serves as the operating system and hosts a diverse range of AI Agent applications, as shown in Figure 1. The LLM as OS (LLMAO) environment shall also provide a comprehensive set of AIOS SDKs and/or libraries, predominately supporting programming in natural language, to help developers or even average users without any knowledge on professional programming languages, to easily develop and deploy Agent applications in the LLMAO-based AIOS-Agent ecosystem.
@@ 2.1.4 Evolution History of Operating Systems
  • Visualization (GUI). As described previously in section 2.1.2, the command-line interface used to be the narrow bridge between users to interact with OS. Since command lines are highly professional, it prevented broader groups of users from easily and efficiently operating the computers. From Xerox Alto introduced by Palo Alto Research Center (PARC) in 1973121212https://spectrum.ieee.org/xerox-alto, to Apple Macintosh introduced in 1984131313http://apple-history.com/128k, to Microsoft Windows [^14] https://winworldpc.com/product/windows-3/31 introduced in the early 1990s, and to a wide range of open-source GUIs for Linux such as GNOME [^15] https://www.gnome.org/, KDE [^16] https://www.kde.org/ and UNITY [^17] https://unityd.org/, the development of GUIs has significantly influenced the accessibility and usability of computers, making computing more intuitive and user-friendly.
@@
2.2 LLMAO, AIOS and AI Agents
@@ 2.3 Development of OS and AIOS Aligned
Figure 3: Parallel Development of OS and AIOS/LLMAO. [^23] Image Sources:
@@
Figure 23 draws a parallel between the evolution of Operating Systems (OS) and LLM as OS (LLMAO). This comparison highlights their parallel progression in terms of enhanced functionality and user interaction. Initially, operating systems such as Atlas Supervisor were designed to manage basic computer resources such as CPU and DRAM. This evolved into more sophisticated version, exemplified by UNIX which can manage various peripheral devices. In a similar vein, LLMs have progressed from text-in-text-out chatbots to intricate LLM-based agents capable of managing various tools for complex task solving, as seen in OpenAGI (Ge et al., 2023). The figure also emphasizes the pivotal role of ARPANET in the development of TCP/IP protocols, which laid the foundation for today’s Internet and connects multiple computers for communicating with each other. This is paralleled by the advancement of LLM-based multi-agent systems where agents can communicate with each other, signifying a trend towards more interconnected and collaborative LLM-based computing environments (Park et al., 2023). Additionally, the evolution of operating systems is marked by the development of sophisticated graphical user interfaces (GUIs), such as those in Windows and Apple Macintosh. Similarly, LLM is evolving to include multi-modal interfaces, as demonstrated by GPT-4V(ision) (OpenAI, 2023). This comparison underscores the role of both OS and LLM in revolutionizing user interaction with technology, transitioning from managing fundamental components to facilitating more intuitive, user-centric experiences.
@@ 3 Architecture of AIOS
In this section, we establish the conceptual framework of “LLM as OS (LLMAO),” creating parallels between LLMAO components and traditional OS elements, as is shown in Table 1. In this analogy, the LLM is compared to the OS kernel, the context window to memory, and external storage to the file system. Tools and user prompts are equated with devices/libraries and the user interface, respectively. We will delve into the specifics of this correlation in the following discussion.
@@ 3.1.1 Reasoning and Planning
In LLMAO, the role of LLM as the kernel subtly echoes the dynamics illustrated in the Dining Philosophers problem from computer science, a scenario that highlights the challenges of resource allocation and synchronization. In this problem, philosophers seated around a table must navigate shared resource usage–forks–in a way that avoids deadlock, a situation where no progress is made due to mutual blocking. Reflecting these challenges, the kernel in an operating system is adept at managing and scheduling resources to avert system deadlocks, ensuring smooth and conflict-free operations. This essential function of the kernel in maintaining system stability and efficiency finds a parallel in the role of the LLM within the AIOS framework. Here, the LLM is responsible for navigating complex informational environments and managing external resources. Crucial to the LLM’s role is its planning ability, which involves generating a series of actions to achieve a specific goal (Ge et al., 2023). Central to the planning ability is the capability of reasoning (Bratman et al., 1988; Russell and Norvig, 1995; Fainstein and DeFilippis, 2015; Huang and Chang, 2022). Through reasoning, LLMAO deconstructs complex tasks from user instructions into more manageable sub-tasks, devising appropriate plans for each. Next, we will explore several representative planning strategies that illustrate this capability.
@@ 3.1.2 Self-Improving
Just as kernel updates are driven by human feedback, focusing on bug reports and performance issues, leading to improvements in functionality, security, and stability, LLMAO also requires continuous enhancement to elevate its performance. This process involves the LLMs learning from interactions, refining their algorithms based on user experiences and feedback. By doing so, LLMs can develop new capabilities and skills, adapting to changing requirements and expectations. This iterative process of improvement ensures that LLMs remain effective, relevant, and efficient in handling diverse tasks and queries, mirroring the evolving nature of technology and user needs. After the LLM has been pre-trained, the learning strategies of LLM can be broadly categorized into two main types, as summarized and exemplified in the following.
@@ 3.2 Context Window (as Memory)
In LLMAO, memory functions similarly to the context window in an LLM, defining the scope of information that the LLM can utilize and learn from while producing outputs. Increasing the amount of memory is desirable but always at a high cost.
@@
  • Position encoding: Position encoding can be classified into absolute position encoding and relative position encoding. The original Transformers paper (Vaswani et al., 2017) utilizes absolute position encoding, in which information is encoded in a combination of sin/cos functions whose wavelength increases from low to higher order dimensions of the embedding. To emphasize the relative position information between tokens, various methods, such as RoPE (Su et al., 2021), Alibi (Sun et al., 2022), and XPos (Press et al., 2021), have been proposed. To increase context length, methods such as position interpolation (Chen et al., 2023b) and NTK-aware (Rozière et al., 2023) position interpolation based on RoPE can be used.
@@ 3.3.2 Data Retrieval Methods
The ability to access relevant and accurate information from external storage through sophisticated data retrieval methods is crucial for LLMAO’s execution of targeted actions. A primary challenge in this process is selecting the most appropriate data file from an extensive repository of information. These data retrieval methods operate automatically, leveraging advanced algorithms and machine learning techniques.
@@ 3.4.1 Tool Categories
  • Self-made Tools. Current tools are primarily designed by humans. Recently, there are increasing interest in the use of LLMs for automated tool making. This involves generating executable programs or enhancing existing tools to create more powerful solutions, guided by appropriate instructions and demonstrations (Qin et al., 2023a; Qian et al., 2023b; Chen et al., 2021). For example, a large code model, as evaluated in Chen et al. (2021), is capable of generating executable programs based on user-provided instructions. These programs can then serve as specialized tools to address particular tasks. Furthermore, these LLMs can also acquire the ability to self-debug, an essential skill for maintaining and improving tool functionality, as detailed in Chen et al. (2023a).

### 09 Dec 2023 v2

# LLM as OS, Agents as Apps: Envisioning AIOS, Agents and the AIOS-Agent Ecosystem

@@ Abstract
In this paper, we begin by introducing the architecture and historical evolution of traditional Operating Systems (OS). Then we formalize a conceptual framework for AIOS through “LLM as OS (LLMOS),” [^1] For convenience, LLMOS may be pronounced as “el-mos”. drawing analogies between AIOS components and traditional OS elements: LLM is likened to OS kernel, context window to memory, external storage to file system, hardware tools to peripheral devices, software tools to programming libraries, and user prompts to user commands. Subsequently, we introduce the new AIOS-Agent Ecosystem, where users and developers can easily program Agent Applications (AAPs) using natural language, democratizing the development of and the access to computer software, which is different from the traditional OS-APP ecosystem, where desktop or mobile applications (APPs) have to be programmed by well-trained software developers using professional programming languages. Following this, we explore the diverse scope of Agent Applications. These agents can autonomously perform diverse tasks, showcasing intelligent task-solving ability in various scenarios. We delve into both single agent systems and multi-agent systems, as well as human-agent interaction. Lastly, we posit that the AIOS-Agent ecosystem can gain invaluable insights from the development trajectory of the traditional OS-APP ecosystem. Drawing on these insights, we propose a strategic roadmap for the evolution of the AIOS-Agent ecosystem. This roadmap is designed to guide the future research and development, suggesting systematic progresses of AIOS and its Agent applications.
@@ 1 Introduction
(b) Architecture of Large Language Model as OS (LLMOS) for AIOS.
Figure 2: Illustrations of the architectures of OS and AIOS (LLMOS).
@@
Inspired by the architecture of traditional OS (as shown in Figure. 1(a)), we present a general framework for LLM as OS (LLMOS) with several key components in Section 3 (as shown in Figure. 1(b)): LLM as Kernel, Context Window as Memory, External Storage as File, Tools as Devices/Libraries, User Prompt/Instruction as User Interface (UI), and Agents as Applications. The conceptual framework presented draws an analogy between an LLM as OS (LLMOS) and a traditional Operating System (OS), mapping various components of the LLMOS to elements of an OS. The LLM itself is likened to the kernel, the central part managing the system’s core functions. The LLM’s context window is compared to the memory of an OS, handling immediate context selection and data processing (Shi et al., 2023). External storage for the LLM is analogous to the files of an OS, allowing for long-term data storage. Meanwhile, there are corresponding data retrieval methods to enable retrieval-augmented LLMs (Guu et al., 2020), which serve as the file system of OS to manage and find relevant files. Besides, LLM can make use of various tools for task solving (Ge et al., 2023), including both hardware tools and software tools. Hardware tools in the LLMOS framework are equated to peripheral devices in a traditional OS, each offering specific functionalities to help interact with the physical world, while software tools in the LLMOS framework serve as the programming libraries in traditional OS, enabling Agent applications to interact with the virtual/digital world. User prompts or instructions for the LLM are akin to the user interface (UI) in a traditional OS, facilitating interaction between the user and the system. The user prompts or instructions can be direct natural language instructions provided by the user, and they may also be (sometimes semi-structured) natural language instructions converted from users’ non-natural language instructions such as clicking on icons. This mapping provides a systematic way to understand the operational similarities between LLMOS-based AIOS and traditional OS.
@@
Upon establishing a robust conceptual framework for LLMOS-based AIOS, we introduce the AIOS-Agent Ecosystem, akin to the traditional OS-APP Ecosystem, in Section 4. We begin by introducing the concept of Agent Applications (AAPs) within LLMOS, analogous to traditional applications (APPs) based on an operating system. These AAPs represent a diverse scope of specialized tasks for users to execute based on LLMOS. As illustrated in Figure 4, by integrating the LLMOS layer with the OS layer, Hardware layer, and the Agent Application layer, we can construct an autonomous AI Agent system. Such AI Agent system responds to user prompts or instructions in natural language and is capable of performing a multitude of tasks based on its interaction with the physical or digital environment. Since a diverse scope of Agents can be developed on top of the shared AIOS foundation, this eventually leads to an AIOS-Agent ecosystem. Moreover, in the new AIOS-Agent Ecosystem, the users and developers can easily program Agent Applications (AAPs) using natural language, democratizing the development of and the access to computer software, which is different from the traditional OS-APP ecosystem, where desktop or mobile applications (APPs) have to be programmed by well-trained software developers using professional programming languages.
@@
Following this, we further delve into the practical LLMOS-based Agent Applications in Section 5. This section explores the potential of enhancing LLMOS’s functionality by developing various agents in the real world, and further investigates the dynamic interactions between multiple agents and humans. LLMOS-based agents are characterized by their creative autonomy, enabling them to generate novel ideas, narratives, or solutions not pre-programmed into them (Chase, 2022; Gravitas, 2023; Ge et al., 2023; Li et al., 2023; Yao et al., 2022b, a), which is indicative of an advanced level of creative intelligence (Yuan et al., 2022; Franceschelli and Musolesi, 2023). Specifically, Section 5.1 discusses applications of single agents, Section 5.2 explores multi-agent systems, and Section 5.3 focuses on human-agent interactions.
@@
Finally, in Section 6, we explore several crucial future research directions for LLMOS and AIOS in general, drawing parallels and learning from the evolution of traditional operating systems. These directions span a wide array of areas, aiming to enhance the capabilities and applications of LLMOS: 1) Resource Management. For example, OS employs virtual and shared memories to address the issue of limited physical memory. LLMOS can inspire from these ideas to mitigate its own problem of limited context window challenges; 2) Communication. Different OSes and applications communicate using standardized protocols (such as Domain-Specific Languages); LLMOSes and Agents can build and use similar standard protocols for exchanging data and instructions with various systems, ensuring compatibility and smooth interaction across diverse platforms; 3) Security. Security vulnerabilities in OS are important issues. State-of-the-art approaches aim to detect and capture malware and viruses at various levels. Similarly, LLMOS can implement detection and intervention mechanisms to regulate and monitor the implementation of third-party tools and Agent Applications.
@@ 2.1.3 Operating System Ecosystem
Similarly, we envision an AIOS-Agent ecosystem, where LLM serves as the operating system and hosts a diverse range of AI Agent applications, as shown in Figure 1. The LLM as OS (LLMOS) environment shall also provide a comprehensive set of AIOS SDKs and/or libraries, predominately supporting programming in natural language, to help developers or even average users without any knowledge on professional programming languages, to easily develop and deploy Agent applications in the LLMOS-based AIOS-Agent ecosystem.
@@ 2.1.4 Evolution History of Operating Systems
  • Visualization (GUI). As described previously in section 2.1.2, the command-line interface used to be the narrow bridge between users to interact with OS. Since command lines are highly professional, it prevented broader groups of users from easily and efficiently operating the computers. From Xerox Alto introduced by Palo Alto Research Center (PARC) in 1973131313https://spectrum.ieee.org/xerox-alto, to Apple Macintosh introduced in 1984141414http://apple-history.com/128k, to Microsoft Windows [^15] https://winworldpc.com/product/windows-3/31 introduced in the early 1990s, and to a wide range of open-source GUIs for Linux such as GNOME [^16] https://www.gnome.org/, KDE [^17] https://www.kde.org/ and UNITY [^18] https://unityd.org/, the development of GUIs has significantly influenced the accessibility and usability of computers, making computing more intuitive and user-friendly.
@@
2.2 AIOS, LLMOS and AI Agents
@@ 2.3 Development of OS and AIOS Aligned
Figure 3: Parallel Development of OS and AIOS/LLMOS. [^24] Image Sources:
@@
Figure 24 draws a parallel between the evolution of Operating Systems (OS) and LLM as OS (LLMOS). This comparison highlights their parallel progression in terms of enhanced functionality and user interaction. Initially, operating systems such as Atlas Supervisor were designed to manage basic computer resources such as CPU and DRAM. This evolved into more sophisticated version, exemplified by UNIX which can manage various peripheral devices. In a similar vein, LLMs have progressed from text-in-text-out chatbots to intricate LLM-based agents capable of managing various tools for complex task solving, as seen in OpenAGI (Ge et al., 2023). The figure also emphasizes the pivotal role of ARPANET in the development of TCP/IP protocols, which laid the foundation for today’s Internet and connects multiple computers for communicating with each other. This is paralleled by the advancement of LLM-based multi-agent systems where agents can communicate with each other, signifying a trend towards more interconnected and collaborative LLM-based computing environments (Park et al., 2023). Additionally, the evolution of operating systems is marked by the development of sophisticated graphical user interfaces (GUIs), such as those in Windows and Apple Macintosh. Similarly, LLM is evolving to include multi-modal interfaces, as demonstrated by GPT-4V(ision) (OpenAI, 2023). This comparison underscores the role of both OS and LLM in revolutionizing user interaction with technology, transitioning from managing fundamental components to facilitating more intuitive, user-centric experiences.
@@ 3 Architecture of AIOS
In this section, we establish the conceptual framework of “LLM as OS (LLMOS),” creating parallels between LLMOS components and traditional OS elements, as is shown in Table 1. In this analogy, the LLM is compared to the OS kernel, the context window to memory, and external storage to the file system. Tools and user prompts are equated with devices/libraries and the user interface, respectively. We will delve into the specifics of this correlation in the following discussion.
@@ 3.1.1 Reasoning and Planning
In LLMOS, the role of LLM as the kernel subtly echoes the dynamics illustrated in the Dining Philosophers problem from computer science, a scenario that highlights the challenges of resource allocation and synchronization. In this problem, philosophers seated around a table must navigate shared resource usage–forks–in a way that avoids deadlock, a situation where no progress is made due to mutual blocking. Reflecting these challenges, the kernel in an operating system is adept at managing and scheduling resources to avert system deadlocks, ensuring smooth and conflict-free operations. This essential function of the kernel in maintaining system stability and efficiency finds a parallel in the role of the LLM within the AIOS framework. Here, the LLM is responsible for navigating complex informational environments and managing external resources. Crucial to the LLM’s role is its planning ability, which involves generating a series of actions to achieve a specific goal (Ge et al., 2023). Central to the planning ability is the capability of reasoning (Bratman et al., 1988; Russell and Norvig, 1995; Fainstein and DeFilippis, 2015; Huang and Chang, 2022). Through reasoning, LLMOS deconstructs complex tasks from user instructions into more manageable sub-tasks, devising appropriate plans for each. Next, we will explore several representative planning strategies that illustrate this capability.
@@ 3.1.2 Self-Improving
Just as kernel updates are driven by human feedback, focusing on bug reports and performance issues, leading to improvements in functionality, security, and stability, LLMOS also requires continuous enhancement to elevate its performance. This process involves the LLMs learning from interactions, refining their algorithms based on user experiences and feedback. By doing so, LLMs can develop new capabilities and skills, adapting to changing requirements and expectations. This iterative process of improvement ensures that LLMs remain effective, relevant, and efficient in handling diverse tasks and queries, mirroring the evolving nature of technology and user needs. After the LLM has been pre-trained, the learning strategies of LLM can be broadly categorized into two main types, as summarized and exemplified in the following.
@@ 3.2 Context Window (as Memory)
In LLMOS, memory functions similarly to the context window in an LLM, defining the scope of information that the LLM can utilize and learn from while producing outputs. Increasing the amount of memory is desirable but always at a high cost.
@@
  • Position encoding: Position encoding can be classified into absolute position encoding and relative position encoding. The original Transformers paper (Vaswani et al., 2017) utilizes absolute position encoding, in which information is encoded in a combination of sin/cos functions whose wavelength increases from low- to higher-order dimensions of the embedding. Various methods, such as RoPE (Su et al., 2021), are proposed later to emphasize the relative position information between tokens. To extend the context window size, Alibi (Sun et al., 2022), LeX (Sun et al., 2022), and XPos (Press et al., 2021) are proposed to enable length extrapolation so that a model can conduct inference on long context while being trained only on short context. To increase context length, methods such as position interpolation (Chen et al., 2023b) and NTK-aware (Rozière et al., 2023) position interpolation based on RoPE can be used.
@@ 3.3.2 Data Retrieval Methods
The ability to access relevant and accurate information from external storage through sophisticated data retrieval methods is crucial for LLMOS’s execution of targeted actions. A primary challenge in this process is selecting the most appropriate data file from an extensive repository of information. These data retrieval methods operate automatically, leveraging advanced algorithms and machine learning techniques.
@@ 3.4.1 Tool Categories
  • Self-made Tools. Current tools are primarily designed by humans. Recently, there are increasing interest in the use of LLMs for automated tool making. This involves generating executable programs or enhancing existing tools to create more powerful solutions, guided by appropriate instructions and demonstrations (Qin et al., 2023a; Qian et al., 2023b; Chen et al., 2021). For example, a large code model, as evaluated in Chen et al. (2021), is capable of generating executable programs based on user-provided instructions. These programs can then serve as specialized tools to address particular tasks. Furthermore, these LLMs can also acquire the ability to self-debug, which is an essential skill for maintaining and improving the tool functionality, as detailed in Chen et al. (2023a).